OpenStudy (mssterehearts):

Factor completely: 5x^2 + 20x Can I have the steps to complete this please?

OpenStudy (skullpatrol):

any ideas?

OpenStudy (mssterehearts):

I don't know what the first step is I think find the GCF

OpenStudy (skullpatrol):

Yes of 5 and 20

OpenStudy (xguardians):

Yes, that is correct, find any values that they both share. x and 5.

OpenStudy (mssterehearts):

They both have GCF of 5

OpenStudy (skullpatrol):

Good, now what is the common factor of x and x^2?

OpenStudy (mssterehearts):

That't where I'm confused at

OpenStudy (skullpatrol):

x^2 means x*x

OpenStudy (mssterehearts):

Is it x lol idk I'm so confused

OpenStudy (skullpatrol):

Yes, exactly.

OpenStudy (mssterehearts):

Yay :D Whats the next step?

OpenStudy (skullpatrol):

Put the two together and you get 5x, right?

OpenStudy (mssterehearts):

Yes

OpenStudy (skullpatrol):

Now factor 5x out of both terms.

OpenStudy (mssterehearts):

Can I have help with that ;/

OpenStudy (skullpatrol):

Recall: ab + ac = a ( b + c )

OpenStudy (skullpatrol):

Here a = 5x

OpenStudy (mssterehearts):

So ab+ac=5x(b+c) ?

OpenStudy (skullpatrol):

5xx + 20x

OpenStudy (mssterehearts):

Oh

OpenStudy (mssterehearts):

I feel dumb because Im not really getting this

OpenStudy (skullpatrol):

20x = 4(5x)

OpenStudy (mssterehearts):

OOhhh okay so you did 20/5 and that how you got 4

OpenStudy (skullpatrol):

Yup

OpenStudy (xguardians):

Imagine dividing both terms by 5. 5(x^2 + 4x) Then divide out an x 5x(x+4)

OpenStudy (mssterehearts):

Yeahh

OpenStudy (xguardians):

So when you multiply these two terms together, you can check your work by seeing you get the same thing you started with.

OpenStudy (mssterehearts):

Makes sense! Thanks!

OpenStudy (mssterehearts):

I needed to understand the steps more than the answer

OpenStudy (skullpatrol):

$$\Huge 5x^2 + 20x = (5x)x + (5x)4$$

OpenStudy (mssterehearts):

Thank you

OpenStudy (skullpatrol):

Thanks for trying to understand the steps :-)